« February 2006 | Main

March 20, 2006

Operations Summary Blog

The operations summary blog has just been updated with the IOP-14 operations summary as well as some final overview plots summarizing the entire HMT-West 2006 project.

Kingsmill

March 07, 2006

END OF PROJECT

After consulting with several participants in the project, it has been decided to end active field operations for HMT-West 2006. Today was the last daily coordination call for this field season. Unattended sensors will continue to operate, although if they malfunction, they will not be repaired. The plan at present is to remove much of the remaining ESRL instrumentation by the middle of April. However, some surface sites, particularly those with soil moisture sensors, will remain in place long term.

As a means to provide an overview of HMT-West 2006 and start to make plans for HMT-West 2007, we would like to hold a workshop in the May-June timeframe, preferably in the Sacramento area. Please provide me your thoughts on this issue (via email or phone). This issue will also be discussed during HMT Advisory Panel conference calls along with other HMT issues of interest. These call will probably begin again in the next month or so.

Many thanks to the forecasters, engineers, instrument operators and coordinators, chief scientists, program managers and other key individuals that made HMT-West 2006 a huge success.

Dave Kingsmill

Project Status: 7 March 2006 (Tuesday)

IOP-14 ended at noon PST. All soundings (OAK & Slough House) have been exhaused. Since IOP-14 began on Sunday 09 UTC Blue Canyon has received 2.51 inches of liquid equivalent precipitation. Only showery precip remains over the ARB and that is expected to continue for the rest of the day then diminish tonight.

In the longer term the next system is a fast-moving one arriving late Wed night into Thursday followed by a series of waves with progressively colder and showery conditions. By the beginning of next week an extended period of drier weather moves in as the mean ridge developes over the West Coast.

--Jorgensen

March 06, 2006

Project Status: 6 March 2006 (Monday)

IOP-14 continues. The precip maximum with the main frontal band was later getting started in the ARB than anticipated yesterday and the precip max was highest just north of the ARB. As of 8 AM PST this morning Blue Canyon had received a 24-hour total of 1.36". The frontal band should pass east of the ARB this afternoon. Cold air advection aloft should bring more convective precip to the area with the higher reaches of the ARB (e.g., Blue Canyon) experiencing near continuous precipitation with snow levels around 4000 feet or so and gradually dropping. The trough axis still looks to pass the ARB around 18 UTC Tuesday morning and more steady precipitation will develop ahead of it beginning around 09 UTC and continuing until the axis passes. An additional 24 hour QPF of 1" is anticipated.

Therefore, IOP-14 will continue at least through noon PST tomorrow.

An announcement on the continuation or termination of HMT field activities will be made on tomorrow's call.

Equipment status: The Alta S-band vertical pointing radar is again working OK as of 5 March 20 UTC. Apparantly a layer of ice on the radome that had been attenuating the signal has melted. The Slough House supplementary soundings ran out at 18 UTC this morning (there were 2 sonde failures last night). The XPOL is working fine and can continue collecting data for at least another 24 hours or so.

--Jorgensen

Project Status: 6 March 2006 (Monday morning)

IOP-14 continues with the x-band polarimetric radar scanning at Auburn and the last of the Slough House soundings being launched this morning. The storm was very tardy getting to the ARB as it stalled east and north of the area. Looks like we began this iop about 15 hours too early. However, steady precip did begin around 00 UTC on 6MAR and continues at sunrise. Alta and Auburn have picked up just over an inch of rain, while Blue Canyon is getting good accumulations of snow.

- Martner

March 05, 2006

Project Status: 5 March 2006 (Sunday)

IOP-14 is undeway. Front is hung up along the coast so precip is light in the ARB. The DAX radar shows heaviest precip north of the ARB. Over 2" has fallen so far north of San Francisco. Trough axis later on Monday will amplify precip a bit before frontal passage and precip begins to wane.

XPOL scanning radar is up and running as are the Slough House supplementary soundings since no reports have been received to the contrary.

Longer range outlook still has another shot of precip Wed-Thurs as a series of short waves impacts the west coast but under much colder conditions.

--Jorgensen

March 04, 2006

Project Status: 4 March 2006 (Saturday)

IOP-14 is scheduled to begin early Sunday morning around 1 AM PST (09 UTC Sunday) with the XPOL scanning radar and Slough House supplementary soundings. Although the forecast is a bit slower arriving in the ARB than yesterday's guidance, the XPOL could begin operations earlier should precip begin before 09 UTC. Heavy precip is expected Sunday evening into Monday. Perhaps 2" of liquid by Monday evening. A frontal passage is expected Monday afternoon then a weakening of the precipiation insensity. Another push of moisture with the trough axis passage Tuesday night through Wednesday may bring another 1.5". A decision on terminating XPOL operations will be made when the precip tapers off late Monday or early Tuesday.

There are 14 Slough House supplementary sondes left and once those are exhaused the HMT Slough House sounding operations are complete. Assuming no sonde failures those sondes, deployed every 3 hours, would run to 00 UTC Tuesday (4 PM PST Monday afternoon) which should cover the fropa.

Sensor Status: The Alta S-band vertical pointing radar is showing evidence of attentuation. Although the antenna cover was brushed off of snow, perhaps some ice buildup is attenuating the beam.

The Blue Canyon precip gage is showing much lower readings that the other gages at the site and is suspect. Other precip sensors (including the ALERT gage) appear to be OK.

A decision on terminating HMT operations after IOP-14 or perhaps after the next precip event later in the week will be made following IOP-14

--Jorgensen

March 03, 2006

Project Status: 3 March 2006 (Friday afternoon)

Precipitation in the American River Basin is presently winding down. IOP-13 ended a short time ago at 20 UTC.

Attention now turns to the next event that is bearing down on northern California. The timing for this event has evolved quite a bit over the last couple forecast cycles, with earlier onsets in evidence for each succeeding forecast. At present, the forecast calls for light precipitation to begin in our target area Saturday night into Sunday morning. Precipitation intensity is forecast to increase during the day on Sunday into Sunday night, with cold frontal passage sometime between 06-12 UTC Monday (Sunday night into Monday morning). Thereafter, the precipitation will linger quite a while, being helped by a few postfrontal cold surges. Precipitation will probably continue into Tuesday. Snow levels will start out around 5-6 kft on Sunday, then descend to 3-4 kft during the Monday-Tuesday time period.

There will be HMT-West operations for this event, which will be IOP-14. Ellen Sukovich will deploy to the area on Saturday to operate XPOL starting Saturday evening. The desired plan for Sloughhouse soundings is to begin operations at 09 UTC Sunday (i.e., Sunday morning) and continue through 00 UTC Tuesday (i.e., Monday afternoon). Dave White is attempting to line up the students to perform these launches. If executed as desired, this plan would exhaust the Sloughhouse sounding resources for the project.

Kingsmill

Project Status: 3 March 2006 (Friday morning)

IOP-13 continues with the X-band polarimetric radar scanning as a wide band-like region of precipitation enters the foothills. Soundings are done for this event. X-band will continue until about noon PST.

- Martner

March 02, 2006

Project Status: 2 March 2006 (Thursday noon)

IOP-13 is in progress with a cold storm that is bringing snow to unusually low elevations (2 kft) in the ARB. Approximately 12 inches of snow fell so far at Blue Canyon where the ALERT gauge shows just over 1 inch of liquid equivalent; similar amounts of liquid equivalent have been recorded at other nearby ALERT gauges and at Foresthill and Alta. This looks like the good cold storm HMT had hoped to get. Soundings at Slough House and supplemental sondes from Oakland continue to be launched from 9 UTC Thursday at 3-hour intervals through 12 UTC Friday. The X-band polarimetric radar began collecting data at 0630 UTC today (10:30 pm PST Wednesday). SMART-R radar was unable to participate because of a transmtter problem.

The x-band will continue at the operator's discretion through the showery daytime weather today, but will definitely be collecting continuous data from this evening and through tomorrow morning, when the precipitation is expected to have a second surge.

The next storm is shaping up to be a fairly good one in the forecasts. It should arrive in the ARB around Sunday afternoon and coud last for 36-48 hours or so into Tuesday. A decision on whether to conduct IOP-14 on that storm will probably be made at tomorrow's conference call. If so, it would involve the X-band radar and the Slough House soundings. The SMART-R radar is now finished for the project and is heading back to Oklahoma. The Oakland supplemental soundings will be finished for the project by tomorow morning.

- Martner

Project Status: 2 March 2006 (Thursday Morning)

IOP-13 began last night at 10 PM PST. Precipitation amounts so far this morning are respectable in the ARB. As of 5 AM Blue Canyon had received 0.63". Snow levels are around 3 kft.

Instrument Status: The SMART-R scanning radar suffered a major transmitter failure soon after midnight. Problem cannot be easily fixed in Foresthill so the radar will depart HMT on Thursday for its return home.

Jorgensen

March 01, 2006

Project Status: 1 March 2006 (Wednesday)

At the present time (20 UTC) partly cloudy skies dominate the ARB. IOP-13 is set up to begin late tonight. Radars should plan on starting operations by 06 UTC (10 PM PST) and continue to at least 12 UTC Friday morning (4 AM PST). The SMART-R will then depart for home, but the XPOL could operate a few hours longer.

Oakland (OAK) supplemental soundings will commence at 09 UTC Thursday morning and end at 09 UTC Friday morning. That will exhause all available NWS supplementary soundings.

Sloughhouse soundings--will commence at 09 UTC Thursday and will continue at 3-hourly intervals to 12 UTC Friday.

The may be a precip lull Thursday mid-day before a second round of heavier rain Thursday night to Friday. Rain should taper off by 18 UTC Friday.

Total QPF for the system(s) may exceed 2.5". Freezing levels will start out around 5 kft and then descend to 2 kft toward the end of the event.

Looking farther ahead there maybe another shot of precip 6-7 March, and again on 10-11 March, but the good tropical feed of PW is drying up beginning now.

--Jorgensen

Project Status: 28 February 2006 (Tuesday)

At the present time (20 UTC), precip over the ARB has moderated considerably from the very wet conditions of the last few days. A few lingering instability showers are bringing spotty precip to the Sierra foothills. Considerable instability still exists at mid-levels. The NWS currently has a tornado warning out for Napa just east of Sacramento. The models indicate that most showers should begin winding down after 06 UTC tomorrow (Wednesday) and sunny skies will, once again, rein over HMT.

However, both GFS and NAM model guidance indicates a short wave will rotate around the strong low off the Pacific Northwest early Thursday morning and bring upslope and baroclinic lift conditions once more to the ARB. So, IOP-13 will begin at 09 UTC Thursday morning (1 AM PST) to catch the expected rain maximum beginning around 12 UTC Thursday. NAM QPF is about 1.5 inches of precip. HPC is also predicting about 2 inches for the period 12 UTC Thursday to 12 UTC Friday.

The current plan for IOP 13:

Scanning radar operations--XPOL and SMART-R should be ready to start operations by 09 UTC 2 Mar. The SMART-R must depart HMT for home no later than 6 AM (PST) Friday, so the latest it could operate is 12 UTC Friday. Should radar echoes linger over the ARB the XPOL could continue to collect data.

Oakland (OAK) supplemental soundings--will commence at 09 UTC 2 Mar and end at 09 UTC 3 Mar (a total of 7 3-hourly supplemental launches. That will exhause all available NWS supplementary soundings.

Sloughhouse soundings--will commence at 09 UTC 2 Mar and will continue at 3-hourly intervals to 12 UTC Friday.

Instrument status issue:

The Colfax (CFC) S-band profiler thought to be repaired is not yet functional. The other S-band profiler at Alta did not get displayed for the last IOP so its status for IOP-13 is uncertain.

Jorgensen